# A1: Asynchronous Test-Time Scaling via Conformal Prediction

### v1

@@ Abstract
Large language models (LLMs) benefit from test-time scaling, but existing methods face significant challenges, including severe synchronization overhead, memory bottlenecks, and latency, especially during speculative decoding with long reasoning chains. We introduce A1 (Asynchronous Test-Time Scaling), a statistically guaranteed adaptive inference framework that addresses these challenges. A1 refines arithmetic intensity to identify synchronization as the dominant bottleneck, proposes an online calibration strategy to enable asynchronous inference, and designs a three-stage rejection sampling pipeline that supports both sequential and parallel scaling. Through experiments on the MATH, AMC23, AIME24, and AIME25 datasets, across various draft–target model families, we demonstrate that A1 achieves a remarkable 56.7x speedup in test-time scaling and a 4.14x improvement in throughput, all while maintaining accurate rejection-rate control, reducing latency and memory overhead, and no accuracy loss compared to using target model scaling alone. These results position A1 as an efficient and principled solution for scalable LLM inference. We have released the code at https://github.com/menik1126/asynchronous-test-time-scaling.
@@ 1 Introduction
Recent advances in large language models (LLMs) drive interest in thinking models—systems that go beyond next-token prediction to emulate human-like reasoning processes (Guo et al., 2025; Muennighoff et al., 2025; McCoy et al., 2024; Shao et al., 2024). These models excel at leveraging complex reasoning chains, especially in test-time scaling settings (Snell et al., 2024; Li et al., 2025; Muennighoff et al., 2025; Zeng et al., 2025).
Test-time scaling (Chen et al., 2025; Muennighoff et al., 2025; Guo et al., 2025) constitutes an entirely new paradigm that enhances the model’s reasoning capabilities by allocating additional computational resources during the inference stage. Typically, test-time scaling can be categorized into two approaches: sequential scaling (Muennighoff et al., 2025; Guo et al., 2025) and parallel scaling (Chen et al., 2025). However, despite its potential, the challenge of efficiently managing increasing input size or complexity during inference remains a critical limitation, hindering the achievement of high-performance deployment.
@@
Parallel scaling, due to its high efficiency in parallel expansion, emerges as a promising paradigm. By parallelly scaling the number of samples, Chen et al. (2025) partially alleviates the high inference latency and memory bottlenecks caused by scaling the token budget reasoning paths, while improving the efficiency of token sampling. Although some methods that adopt confidence-based early stopping of reasoning chains (Huang et al., 2025; Wan et al., 2024) improve parallel sampling efficiency, problems still remain in memory efficiency and high inference latency. Another potential issue is that early stopping prunes away potentially correct reasoning paths and reduces the diversity of the output space.
Speculative decoding (Li et al., 2024a; Leviathan et al., 2023; Kim et al., 2023; Pan et al., 2025b; Yang et al., 2025) represents a promising approach for accelerating test-time scaling. In this framework, a lightweight draft model generates tokens, which are subsequently validated by a target model. This dual-phase approach not only speeds up inference by offloading most of the generation process to the draft model, but also ensures that the final outputs retain high fidelity, thereby achieving a favorable balance between efficiency and accuracy.
However, when speculative decoding (Pan et al., 2025b; Yang et al., 2025) meets test-time scaling, the decoding process faces two key challenges. The first is the prefill stage memory bottleneck of the target model, which can exhaust memory in parallel scaling scenarios. As shown in Figure 1, as the number of sampling increases, the memory overhead of the target model tends to grow due to KV cache accumulation. This effect becomes more pronounced in target models when attempting to scale the number of requests from the draft model. During real-world deployment on the SGLang server (Zheng et al., 2024), high-concurrency sampling can lead to memory peaks that easily exceed the GPU’s maximum capacity, potentially causing the server to crash. Therefore, it is crucial to constrain the request budget from the draft model to the target model within a manageable range.
@@
The second bottleneck is the need for the target model to rank all samples from each draft model’s inference step before deciding which steps to accept. This presents a synchronization time bottleneck. As illustrated in Fig. 2, during the multi-turn sampling process (with a sampling quantity of 3 at each step), if the target model aims to reject the paths with the lowest two confidences, a ranking operation must be performed at each step. Given the limited computational and memory resources, the target model needs to prioritize processing the most important requests. Especially in test-time scaling, when combining sequential and parallel scaling, the synchronization overhead from precise budget control and the pursuit of globally optimal ranking is amplified. Although this issue has been widely discussed in the context of tool calls (Gim et al., 2024; Ginart et al., 2024), it has not been formally proposed in the context of test-time scaling.
To analyze the synchronization bottleneck and address the two challenges mentioned above, we first introduce a novel variant of arithmetic intensity called asynchronous arithmetic intensity to analyze the system bottleneck, and then explore conformal prediction (Vovk et al., 2005; 2003; Romano et al., 2020; Lei et al., 2018) for prediction ranking to design the asynchronous algorithm. In our formulation, conformal prediction defines a prediction set $C_{\alpha}$; sampling in $C_{\alpha}$ are rejected, while sampling outside are accepted. This yields a distribution-free guarantee that the right sampling is retained (i.e., lies outside $C_{\alpha}$) with high probability, enabling asynchronous test-time scaling. This paper presents the following contributions:
  • We propose asynchronous arithmetic intensity, a new metric designed to characterize and quantify throughput/latency bottlenecks that emerge in test-time scaling scenarios.
  • We introduce conformal prediction to address the memory bottleneck risks and the challenge of prediction ranking.
  • We propose A1 and implement a training-free, lossless acceleration, achieving a 56.7x speedup in test-time scaling and a 4.14x improvement in throughput through sequential and parallel test-time scaling.
@@ 2 Preliminary
In this section, we first introduce how to construct the prediction set in classic setup, and then present our setup.
@@
denote the calibration dataset. Each pair $(X_{i},Y_{i})$ for $i=1,\dots,n$ is a data point, consisting of an $X_{i}$ (the input question for the $i$-th example) and a ground-truth denoted as $Y_{i}$. The symbol $n$ represents the total number of calibration dataset.
$(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m}).$ (2)
In adaptive conformal prediction (Romano et al., 2020; Angelopoulos et al., 2020; Huang et al., 2023), the conformity score for each candidate path is computed using a softmax function:
@@
$\tau=Q_{p}\left(\left\{s_{i}^{j}\mid i=1,\dots,n;\ j=1,\dots,m\right\}\right)$ (4)
where
@@
and $\alpha\in(0,1)$ is the user-specified miscoverage rate. The quantile $Q_{p}$ returns the $p$-th percentile smallest score, ensuring that the resulting prediction set includes the ground-truth with probability at least $1-\alpha$.
@@
In the asynchronous test-time scaling setup, we leverage a draft model for fast sampling and delegate verification to a slower target model. Unlike classical rejection sampling (Chen et al., 2023), the target model evaluates the sampled chains of thought and proactively initiates its own resampling whenever a candidate is rejected, rather than passively awaiting new inputs. This rejection sampling process is repeated until the maximum number of scaling turns is reached, which inevitably introduces additional overhead in this setting. To mitigate synchronization and memory overhead on the target model’s server, we aim to predict and control the rejection rate by constructing prediction sets with bounded size. Conformal prediction offers a principled way to achieve this. Given a predefined $\alpha$, we estimate a confidence level such that the ground-truth $y$ falls within the prediction set $C_{\alpha}(Y)$ with probability at least $1-\alpha$:
$\vskip-5.69054pt\mathbb{P}(y\in C_{\alpha}(Y))\geq 1-\alpha,$ (6)
where $\alpha$ is conventionally interpreted as the significance level (e.g., 0.05 corresponding to 95% confidence). In this work, however, we reinterpret $\alpha$ as the rejection rate of the target model.
@@
However, employing conformal prediction also suffers from the effect of synchronization (Zhu et al., 2025). Specifically, in conformal prediction, the conformal scores assigned to the measured objects are typically required to be normalized, as shown in Eq.3, and this normalization process inherently introduces a bottleneck to parallelization. Moreover, in typical inference engines (Zheng et al., 2024; Kwon et al., 2023), particularly those with asynchronous scheduling, obtaining the normalized scores for all sampling in different batches can be challenging. To address these challenges, we reformulate the task of constructing prediction sets as a ordinal classification (Dey et al., 2023; Xu et al., 2023), where the size of the prediction set corresponds to a reject rate of $\alpha$. Formally, we aim to ensure:
@@
that is, to provide a probabilistic guarantee for the sampled outputs corresponding to each input instance. Our setup focuses on developing asynchronous algorithms for ordinal prediction, with the construction of the set ensuring that its size exactly matches the predefined size under both marginal and conditional coverage.
@@ 3 Challenge and Design
To identify performance bottlenecks in classic setup, we focus on the arithmetic intensity (Spector & Re, 2023), which quantifies the utilization of arithmetic units. Arithmetic intensity is defined as:
@@
where $F$ denotes the number of floating-point operations (FLOPs), and $B$ denotes the number of bytes accessed.
@@ 3.1 Q1: What are the Emerging Performance Bottlenecks?
Speculative decoding (Leviathan et al., 2023) accelerates inference by overlapping computation with speculative memory accesses, enabling multiple draft tokens to be validated in parallel. Its main bottleneck is parallel score computation (Yin et al., 2024), making the process computation-bound.
Building upon this perspective, asynchronous scaling can be seen as an even more aggressive parallelization strategy. The target model validates far more tokens in parallel than in speculative decoding, which intensifies the prefill bottleneck and results in total computation time far exceeding total memory access time, as illustrated by the comparison between the green and yellow lines in Fig. 4(a). In addition, synchronization delays caused by varying sampling speeds across draft steps directly degrade computational resource utilization, which is not captured by Eq. 9. As shown in Fig. 3(b), synchronization overhead grows exponentially with the number of sampling turns. In the parallel scaling setting (Fig. 3(a)), this overhead increases linearly with the number of concurrent paths.
To this end, we observe Fig. 4(a) that increasing the sampling size naturally raises arithmetic intensity (with memory access time being negligible). However, due to synchronization overhead, the overall inference time of the system is increasing. To account for synchronization costs within arithmetic intensity, we define an asynchronous arithmetic intensity $r$:
@@
where $T_{c}$ is computation time, $T_{m}$ is memory access time, $t_{c}$ and $t_{m}$ are the per-unit costs of computation and memory access, respectively. It can be observed from Fig. 4(b) that under synchronization settings, $r$ decreases as the sampling size increases which indicates that, in test-time scaling supported by speculative decoding, synchronization overhead emerges as the primary bottleneck.
@@ 3.2 Q2: How is the prediction set constructed?
Adaptive conformal prediction typically relies on a held-out calibration set to determine the threshold $\tau$, which ensures coverage guarantees. However, in the test-time scaling setup, held-out examples are generally unavailable during inference, making it impractical to reserve a dedicated calibration split. To address this limitation, we propose an online calibration strategy. Specifically, $m$ outputs are pre-sampled for each input in the test set, yielding $(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m})$. We then relax the strict requirement that conformity scores sum to one, and instead directly define:
$s_{i}^{k}=-\ell(X_{i},\hat{Y}_{i}^{k}).$ (11)
This formulation is used to estimate marginal conformal p-values for rejection sampling:
@@
In this formulation, $s_{\xi}^{k}$ denotes the conformity score of the test-time candidate $\hat{Y}_{\xi}^{k}$, which represents the $k$-th sample of the $\xi$-th input on the test set, and $s_{i}^{j}$ are the scores from the calibration set $(X_{i},\hat{Y}_{i}^{j})$. The indicator function $\mathbf{1}(\cdot)$ returns 1 when the condition is satisfied. The resulting conformal $p$-value governs rejection sampling: a candidate is accepted if $p_{n+1}^{k}>\alpha$, ensuring that only high-confidence outputs are retained.
This procedure guarantees marginal coverage at level $1-\alpha$ by selecting the $\lceil p\cdot n\cdot m\rceil$-th smallest score. Conditional coverage can be achieved by adjusting the quantile threshold by selecting the $\lceil p\cdot m\rceil$-th smallest score. A detailed comparison of these two approaches is provided in Appendix A.4.
@@
Let $B$ denote the predefined budget (i.e., the number of candidates to reject). Given a test-time input $X_{i}$, we sample $m$ candidate CoTs $\big(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m}\big)$. For simplicity, we will not distinguish between conditional coverage and marginal coverage in what follows, and mweishenmweishen represents the total number of samples; we then compute their corresponding $p$-values $p_{i}^{1},\dots,p_{i}^{m}$}.
Importantly, this sampling and evaluation process is conducted asynchronously: each candidate $\hat{Y}_{i}^{k}$ is generated independently and evaluated for its $p$-value without requiring synchronization with other candidates. As a result, the outputs implicitly exhibit a descending order:
@@
This formulation ensures that the selected candidates meet the conformal coverage criterion and can be safely retained. Equivalently, this can be interpreted as rejecting the top-$B$ lowest-$p$ candidates.

### v2

# ATTS: Asynchronous Test-Time Scaling via Conformal Prediction

@@ Abstract
Large language models (LLMs) benefit from test-time scaling but are often hampered by high inference latency. Speculative decoding is a natural way to accelerate the scaling process; however, scaling along both the parallel and sequential dimensions poses significant challenges, including substantial memory-bound execution and synchronization overhead. We introduce ATTS (Asynchronous Test-Time Scaling), a statistically guaranteed adaptive scaling framework that follows the hypothesis testing process to address these challenges. By revisiting arithmetic intensity, ATTS identifies synchronization as the primary bottleneck. It enables asynchronous inference through online calibration and proposes an ordinal classification algorithm that supports a three-stage rejection sampling pipeline, scaling along both the sequential and parallel axes. Across experiments on the MATH, AMC23, AIME24, and AIME25 datasets and across multiple draft–target model families, we show that ATTS delivers up to 56.7x speedup in test-time scaling and a 4.14x throughput improvement, while maintaining accurate control of the rejection rate, reducing latency and memory overhead, and incurring no accuracy loss. By scaling both in parallel and sequential dimensions, we enable the 1.5B/70B draft/target model combination to achieve the performance of the state-of-the-art reasoning model o3-mini (high) on the AIME dataset. We have released the code at https://github.com/menik1126/asynchronous-test-time-scaling.
@@ 1 Introduction
With the rapid advances in large language models (LLMs), attention is increasingly turning to reasoning models (Guo et al., 2025; Muennighoff et al., 2025; McCoy et al., 2024; Shao et al., 2024)—systems that transcend next-token prediction in order to emulate human-like reasoning behaviors. These models excel at leveraging complex reasoning chains, especially in test-time scaling settings (Snell et al., 2024; Li et al., 2025; Muennighoff et al., 2025; Zeng et al., 2025).
Test-time scaling (Chen et al., 2025; Muennighoff et al., 2025; Guo et al., 2025) constitutes a new paradigm that enhances the model’s reasoning capabilities by allocating additional computational resources during the inference stage. Typically, test-time scaling can be categorized into two approaches: sequential scaling (Muennighoff et al., 2025; Guo et al., 2025) and parallel scaling (Chen et al., 2025). However, despite its potential, the challenge of efficiently managing increasing sampling size or complexity during inference remains a critical limitation, hindering the achievement of high-performance deployment.
Benefiting from the shared-prefix mechanism of the inference engines (Kwon et al., 2023; Zheng et al., 2024), parallel scaling (Chen et al., 2025) increases the number of samples concurrently, thereby partially mitigating the inference-time latency and memory footprint introduced by scaling the per-trajectory token budget (i.e., longer reasoning paths), while simultaneously improving token-sampling throughput.
Although some methods (Huang et al., 2025; Wan et al., 2024) that adopt confidence-based early stopping of reasoning chains improve parallel sampling efficiency, problems still remain in memory efficiency and high inference latency. Another potential issue is that early stopping prunes away potentially correct reasoning paths and reduces the diversity of the output space.
@@
Speculative decoding (Li et al., 2024a; Leviathan et al., 2023; Kim et al., 2023; Pan et al., 2025b; Yang et al., 2025) represents a promising approach for accelerating decoding. In this framework, a lightweight draft model generates tokens, which are subsequently validated and refined by a target model. This dual-phase approach not only speeds up inference by offloading most of the generation process to the draft model, but also ensures that the final outputs retain high fidelity, thereby achieving a favorable balance between efficiency and accuracy.
However, when speculative decoding (Pan et al., 2025b; Yang et al., 2025) meets test-time scaling, the decoding process faces two key challenges. The first is the memory bottleneck of the target model during the prefill phase. As shown in Figure 1, as the number of sampling increases, the memory overhead of the target model tends to grow due to KV cache accumulation. This effect becomes more pronounced in target models when attempting to scale the number of requests from the draft model. During real-world deployment on the SGLang server (Zheng et al., 2024), high-concurrency sampling, especially when simultaneously validating multiple long reasoning chains, can lead to memory peaks that easily exceed the GPU’s maximum capacity, causing the server to crash. Therefore, it is crucial to constrain the request budget from the draft model to the target model within a manageable range.
@@
In speculative decoding, one stage involves rejection sampling: prior to acceptance, the target model either ranks draft-generated candidates or computes a divergence between the draft and target distributions, introducing an additional synchronization overhead bottleneck. As illustrated in Fig. 2, during the multi-turn sampling process (with a sampling quantity of 3 at each turn), if the target model aims to reject the sampling with the lowest two confidences, a ranking operation must be performed at each turn. Given the limited computational and memory resources, the target model needs to prioritize processing the most important requests. Especially in test-time scaling, when combining sequential and parallel scaling, the synchronization overhead from precise budget control and the pursuit of globally optimal ranking is amplified. Although this issue has been widely discussed in the context of tool calls (Gim et al., 2024; Ginart et al., 2024), it has not been formally proposed in the context of test-time scaling.
To analyze the synchronization bottleneck and address the two challenges mentioned above, we first introduce a novel variant of arithmetic intensity called asynchronous arithmetic intensity to analyze the system bottleneck, and then explore conformal prediction (Vovk et al., 2005; 2003; Romano et al., 2020; Lei et al., 2018) for ranking predictions to design the asynchronous algorithm. In our formulation, conformal prediction defines a prediction set $C_{\alpha}$; sampling in $C_{\alpha}$ are rejected, while sampling outside are accepted. This yields a distribution-free guarantee that the right sampling is retained (i.e., lies outside $C_{\alpha}$) with high probability, enabling asynchronous test-time scaling. This paper presents the following contributions:
  • We propose asynchronous arithmetic intensity, a performance metric designed to characterize and quantify throughput/latency bottlenecks that emerge in test-time scaling scenarios.
  • We introduce conformal prediction to tackle prediction ranking, and—leveraging the resulting ranking—construct stable prediction sets that mitigate GPU-memory bottleneck risks.
  • We propose ATTS, a training-free, lossless acceleration method that achieves a 56.7x speedup in test-time scaling and a 4.14x throughput improvement in both sequential and parallel settings.
@@ 2 Preliminary
We first introduce how to build the prediction set in the classical setup, and then present our setup.
@@
denote the calibration dataset. Each pair $(X_{i},Y_{i})$ for $i=1,\dots,n$ is a data point, consisting of an $X_{i}$ (the input question for the $i$-th example) and a ground truth denoted as $Y_{i}$. The symbol $n$ denotes the size of the calibration dataset. For each input $X_{i}$, we draw $m$ candidate sampling
$(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m}),$ (2)
where $m$ is the number of samples per input (the sampling budget), and $\hat{Y}_{i}^{k}$ denotes the $k$-th candidate sample. In adaptive prediction-set construction (Romano et al., 2020; Angelopoulos et al., 2020; Huang et al., 2023), the conformity score for each sample is computed via a softmax function:
@@
$\tau=Q_{p}\left(\left\{s_{i}^{j},\middle|,i=1,\dots,n,;j=1,\dots,m\right\}\right),$ (4)
@@
and $\alpha\in(0,1)$ is the user-specified miscoverage rate. To satisfy conditional coverage, The prediction set for input $X_{i}$ is then defined as
$C_{\alpha}(X_{i})=\left\{\hat{Y}_{i}^{k},\middle|,s_{i}^{k}\geq\tau,\ k=1,\dots,m\right\}.$ (6)
which ensures that the resulting set includes the ground truth with probability at least $1-\alpha$. Although conformal prediction in the classical setting can provide guaranteed conditional coverage, the conformal scores assigned to candidate outputs are typically required to be normalized, as shown in Eq. 3, and this normalization inherently introduces a bottleneck to parallelization. Therefore, in the subsequent Ordinal Classification, we transform the problem into a hypothesis testing framework via $p$-values to avoid normalization, with a proof of its coverage guarantee provided in Appendix A.7.
@@
In the asynchronous test-time scaling setup, we leverage a draft model for fast sampling and delegate verification to a slower target model. Unlike classical rejection sampling (Chen et al., 2023), which approximates a target distribution with a draft distribution, we focus on accurately predicting the rejection rate, thereby reducing VRAM out-of-memory risk and the synchronization overhead caused by global ranking or softmax function. Given a predefined $\alpha$, we estimate a confidence level such that the ground truth $y$ falls within the prediction set $C_{\alpha}(Y)$ with probability at least $1-\alpha$:
$\mathbb{P}(y\in C_{\alpha}(Y))\geq 1-\alpha,$ (7)
where $\alpha$ is conventionally interpreted as the significance level (e.g., $0.05$ corresponding to $95\%$ confidence). In this work, however, we reinterpret $\alpha$ as the rejection rate of the target model.
@@
In typical inference engines (Zheng et al., 2024; Kwon et al., 2023), particularly those with asynchronous scheduling, obtaining the normalized scores for all sampling in different batches can be challenging. To avoid normalization and global ranking operations, we reformulate the task of constructing prediction sets as an ordinal classification (Dey et al., 2023; Xu et al., 2023), meaning that we predict the ranks of all samples. Formally, we aim to ensure:
@@
That is, it provides a probabilistic guarantee for the sampled outputs corresponding to each input instance. To achieve this, our setup focuses on developing asynchronous algorithms for ranked prediction, where the construction of the prediction set ensures that its size matches the predefined budget while maintaining both marginal and conditional coverage. This approach avoids the need for normalization while addressing the challenges posed by asynchronous scheduling.
@@ 3 Challenge and Design
To identify performance bottlenecks in the classic setup, we introduce arithmetic intensity (Spector & Re, 2023), which measures the utilization of arithmetic units. It is defined as:
@@
where $F$ is the number of floating-point operations (FLOPs) and $B$ is the number of bytes accessed.
@@ 3.1 Q1: What are the Emerging Performance Bottlenecks?
Speculative decoding (Leviathan et al., 2023) accelerates inference by overlapping computation with memory accesses, enabling multiple draft tokens to be validated in parallel. Its main bottleneck is parallel score computation (Yin et al., 2024), making the process computation-bound.
Building upon this perspective, asynchronous scaling can be seen as an even more aggressive parallelization strategy. The target model validates far more tokens in parallel than in speculative decoding, which intensifies the prefill bottleneck and results in total computation time far exceeding total memory access time, as illustrated by the comparison between the green and yellow lines in Fig. 4(a).
As shown in Fig. 3(b), synchronization overhead grows exponentially with the number of sampling turns. In the parallel scaling setting (Fig. 3(a)), this overhead increases linearly with the number of concurrent samples. To this end, we observe Fig. 4(a) that increasing the sampling size naturally raises arithmetic intensity (with memory access time being negligible). To account for synchronization costs within arithmetic intensity, we define an asynchronous arithmetic intensity $r$:
@@
where $T_{c}$ is computation time, $T_{m}$ is memory access time, $t_{c}$ and $t_{m}$ are the per-unit costs of computation and memory access, respectively. It can be observed from Fig. 4(b) that under classic setups, $r$ decreases as the sampling size increases which indicates that synchronization overhead emerges as the primary bottleneck.
@@ 3.2 Q2: How is the prediction set constructed?
Conformal prediction typically relies on a held-out calibration set to determine the threshold $\tau$. However, in the test-time scaling setup, held-out examples are generally unavailable. To address this limitation, we propose an online calibration strategy. Specifically, $m$ outputs are pre-sampled for each input in the test set, yielding $(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m})$. We then relax the strict requirement that conformity scores sum to one, and instead directly define:
$s_{\xi}^{k}=-\ell(X_{\xi},\hat{Y}_{\xi}^{k}).$ (12)
This formulation is used to estimate conformal p-values for rejection sampling:
@@
In this formulation, $s_{\xi}^{k}$ denotes the conformity score of the test-time candidate $\hat{Y}_{\xi}^{k}$, which represents the $k$-th sample of the $\xi$-th input on the test set, and $s_{i}^{j}$ are the scores from the calibration set $(X_{i},\hat{Y}_{i}^{j})$. The indicator function $\mathbf{1}(\cdot)$ returns 1 when the condition is satisfied. The $p$-value based on formula 13 guarantees marginal coverage at level $1-\alpha$, which can be intuitively explained as calculating one’s rank by comparing the conformity score with the score from the $p\cdot n\cdot m$ in the entire calibration set. Conditional coverage can be achieved by adjusting the comparison with the $p\cdot m$ calibration set from the current input sample.
A detailed proof of these two approaches is provided in Appendix A.7. The conformal \$p\$-value governs rejection sampling: a candidate is accepted if $p_{\xi}^{k}>\alpha$, ensuring that only high-confidence outputs are retained, thereby achieving precise budget control.
@@
Let $B$ denote the predefined budget (i.e., the number of candidates to reject). Given a test-time input $X_{i}$, we sample $m$ candidate CoTs $\big(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m}\big)$ in each turn and then compute their corresponding $p$-values $p_{i}^{1},\dots,p_{i}^{m}$}.
Importantly, this sampling and evaluation process is conducted asynchronously: each candidate is generated independently and evaluated for its $p$-value without requiring synchronization with other candidates. As a result, the outputs implicitly exhibit a descending order:
@@
This formulation ensures that the selected candidates meet the coverage rate. Equivalently, this can be interpreted as rejecting the top-$B$ candidate sampling.